OpenStudy (anonymous):

You have worked these hours this week: 5 4/5, 6 1/3, 8 2/5, 4 2/3 how many hours did you work? A. 25 1/5 B. 24 1/5 C. 24 6/5 D. 22 2/5

OpenStudy (misty1212):

add

OpenStudy (misty1212):

use a calculator if necessary

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The denominators you have are 5 and 3. The Least Common Multiple of 5 and 3 is 15, so use 15 as your common denominator. 7 3/5 = 7 9/15 8 2/3 = 8 10/15 ---------------------- sum = (7+8) + (9+10)/15 = 15 + 19/15 = 15 + 1 4/15 = 16 4/15 ::::: 5 4/5 = 5 12/15 6 1/3 = 6 5/15 8 2/5 = 8 6/15 4 2/3 = 4 10/15 ----------------------- sum = (5+6+8+4) + (12+5+6+10)/15 = 23 + 33/15 = 23 + 2 3/15 = 25 1/5

OpenStudy (anonymous):

thanks I have more it just that I ain't good with math and I try to practice but it to hard

OpenStudy (anonymous):

Never mind, keep practicing :) You would get better soon :) Practice makes a man perfect (Hehe, practice also makes a women perfect too :D )

OpenStudy (anonymous):

okay